OpenStudy (anonymous):

−2(x + 3) = −2(x + 1) − 4 2 3 All real numbers No solution

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

start by distributing

OpenStudy (anonymous):

i am like a really slow learner

OpenStudy (anonymous):

a(x+b)=ax + ab

OpenStudy (anonymous):

so i guess the answer is c

OpenStudy (anonymous):

sorry i dont work with guess, u work i help u guess i leave

OpenStudy (anonymous):

ok but i really done understand

OpenStudy (anonymous):

dont

OpenStudy (anonymous):

−2(x + 3) distribute the -2

OpenStudy (anonymous):

−2(x + 3) = −2(x + 1) − 4 Distribute. -2(x + 3) = -2x - 6 -2(x + 1) = -2x - 2 Now you have: -2x - 6 = -2x - 2 - 4 Can you combine like terms?

OpenStudy (anonymous):

yes i do but not in these kind of math problems

OpenStudy (anonymous):

Okay. -2x - 6 = -2x - 2 - 4 What is -2 - 4?

OpenStudy (anonymous):

\[-2x - 6 = -2x (- 2 - 4) \](-2 - 4) = -6 \[-2x - 6 = -2x - 6 \]Add 6 to both sides.\[-2x - 6 + 6= -2x - 6 + 6 \]\[-2x + 0 = -2x + 0\]\[-2x = -2x\]Divide both sides by -2.\[x = x\]